what is f(x)=96+12x+8x+x^2 in vertex form
Answer by stanbon(75887)

f(x)=96+12x+8x+x^2 in vertex form
--------------
x^2 + 20x = y-96
---------
x^2 + 20x+100 = y-96+100
----
(x+10)^2 = y+4
========================
Cheers,
Stan H.
========================